You want 14L of fruit punch that is 10% juice. at the store, you find punch that it 15% juice and punch that is 8% juice. how much of each should you purchase?
Answer by mananth(16949)

-----------------percent----------------quantity
punch Type I 15 ----------------x liters
Punch type II 8 ----------------14-xliters
Mixture 10.00 ---------------- 14

15x+8(14-x)=10*14

15x+112 -8x=140
15x-8x=140-112
7x=28
/7
x=4liters 15.00% punch Type I
10 liters 8.00% Punch type II
m.ananth@hotmail.ca